Publications on Experimental High-Energy Physics at Princeton University

Compiled by K.T. McDonald

(October 22, 2022)

Experimental HEP Personnel are listed at

http://kirkmcd.princeton.edu/hep-group/hep_group.pdf

which includes a list of PhD theses. Scans of these theses are available at

http://kirkmcd.princeton.edu/hep-group/PhD_Theses/

Reynolds Group

1. G.T. Reynolds and W.D. Hardin, Penetrating Power of Extensive Shower Particles,

  Phys. Rev. 74, 1549 (1948), http://kirkmcd.princeton.edu/examples/EP/reynolds_pr_74_1549_48.pdf

2. G.T. Reynolds, On the Effect of Lead above Apparatus for the Detection of Low Energy Muons, Rev. Mod. Phys. 721, 122 (1949),

   http://kirkmcd.princeton.edu/examples/EP/reynolds_rmp_21_122_49.pdf

3. G.T. Reynolds, F.B. Harrison and G. Salvini, Liquid Scintillation Counters, Phys.

  Rev. 78, 488 (1950), http://kirkmcd.princeton.edu/examples/detectors/reynolds_pr_78_488_50.pdf

4. R. Ronald Rau and G.G. Harris, Wide Angle Sprays Produced by Cosmic Radiation,

  Phys. Rev. 79, 915 (1950), http://kirkmcd.princeton.edu/examples/EP/rau_pr_79_915_50.pdf

5. F.B. Harrison and G.T. Reynolds, Spectral Emission from Scintillation Solutions and Crystals, Phys. Rev. 79, 732 (1950),

   http://kirkmcd.princeton.edu/examples/detectors/harrison_pr_79_732_50.pdf

6. R. Ronald Rau, The Nature of Wide-Angle Sprays. I, Phys. Rev. 80, 914 (1950),

7. R. Ronald Rau and A.S. Wightman, The Nature of Wide-Angle Sprays. II, Phys.

  Rev. 80, 914 (1950), http://kirkmcd.princeton.edu/examples/EP/rau_pr_80_914b_50.pdf

8. F.B. Harrison, J.W. Keuffel and G.T. Reynolds, Capture of μ-Mesons in Heavy Elements, Phys. Rev. 83, 680 (1951),

   http://kirkmcd.princeton.edu/examples/EP/harrison_pr_83_680_51.pdf

9. T.N.K. Godfrey, F.B. Harrison and J.W. Keuffel, Satellite Pulses from Photomultipliers, Phys. Rev. 84, 1248 (1951),

   http://kirkmcd.princeton.edu/examples/detectors/godfrey_pr_84_1248_51.pdf

10. J. Keuffel, F.B. Harrison, T.N.K. Godfrey and G.T. Reynolds, Capture of μ-Mesons in Heavy Elements, Phys. Rev. 87, 942 (1952),

   http://kirkmcd.princeton.edu/examples/EP/keuffel_pr_87_942_52.pdf

11. T.N.K. Godfrey, Negative μ-Meson Capture in Carbon, Phys. Rev. 92, 512 (1953),

    http://kirkmcd.princeton.edu/examples/EP/godfrey_pr_92_512_53.pdf

12. G.T. Reynolds and S.B. Treiman, Kinematics of Λ^{0} Production, Phys. Rev. 94, 207

   (1954), http://kirkmcd.princeton.edu/examples/EP/reynolds_pr_94_207_54.pdf

13. L. Mezzetti and J.W. Keuffel, Decay Curve of K Particles, Phys. Rev. 95, 858

   (1954), http://kirkmcd.princeton.edu/examples/EP/mezzetti_pr_95_858_54.pdf

14. A.L. Hodson, J. Ballam, W.H. Arnold, D.R. Harris, R.R. Rau, G.T. Reynolds and S.B. Treiman, Cloud-Chamber Evidence for a Charged Counterpart of the θ^{0} Particle, Phys. Rev. 96, 1089 (1954),

    http://kirkmcd.princeton.edu/examples/EP/hodson_pr_96_1089_54.pdf

15. S.B. Treiman G.T. Reynolds and A.L. Hobson, Angular Correlation Effects in Unstable Particle Decay, Phys. Rev. 97, 244 (1955),

    http://kirkmcd.princeton.edu/examples/EP/treiman_pr_97_244_55.pdf

16. J. Ballam, A.L. Hodson, W. Martin, R.R. Rau, G.T. Reynolds and S.B. Treiman, Orientation of Planes in Double V^{0} Decay Events, Phys. Rev. 97, 245 (1955),

    http://kirkmcd.princeton.edu/examples/EP/ballam_pr_97_245_55.pdf

17. W.H. Arnold, J. Ballam, G.K. Lindberg and V.A.J. van Lint, Example of the Cascade Decay of a Negative Hyperon, Phys. Rev. 98, 838 (1955),

    http://kirkmcd.princeton.edu/examples/EP/arnold_pr_98_838_55.pdf

18. J. Ballam, A.L. Hodson and G.T. Reynolds, Observations on S Particles, Phys. Rev.

   99, 1038 (1955), http://kirkmcd.princeton.edu/examples/EP/ballam_pr_99_1038_55.pdf

19. G.T. Reynolds and W.A. Aron, Masses of S Particles, Phys. Rev. 99, 1038 (1955),

    http://kirkmcd.princeton.edu/examples/EP/reynolds_pr_99_1038_55.pdf

20. W.H. Arnold, J. Ballam and G.T. Reynolds, Cloud-Chamber Observations on Charged V Particles, Phys. Rev. 100, 295 (1955),

    http://kirkmcd.princeton.edu/examples/EP/arnold_pr_100_295_55.pdf

21. W.H. Arnold Jr, W. Martin and H.W. Wyld, Anomalous Neutral V -Particles, Phys.

   Rev. 100, 1545 (1955), http://kirkmcd.princeton.edu/examples/EP/arnold_pr_100_1545_55.pdf

22. T. Bowen, J. Hardy Jr, G.T. Reynolds, G. Tagliaferri, A.E. Werbrouck and W.H. Moore, Observation of Unpolarized Λ’s Produced by 1.5-Bev π^{−} Interactions in Pb, Fe, and C, Phys. Rev. Lett. 1, 11 (1960),

    http://kirkmcd.princeton.edu/examples/EP/bowen_prl_1_11_58.pdf

23. R. Giaconni, W.H. Blum and G.T. Reynolds, Detection of High Energy μ-Mesons by an Air Čerenkov Counter, Nuovo Cim. 11, 102 (1959),

    http://kirkmcd.princeton.edu/examples/detectors/giaconni_nc_11_102_59.pdf

24. G.T. Reynolds, R. Giaconni and D. Scarl, Cosmic-Ray Tracks in Filament Scintillation Chambers, Rev. Sci. Instr. 30, 497 (1959),

http://kirkmcd.princeton.edu/examples/detectors/reynolds_rsi_30_497_59.pdf

25. T. Bowen, J. Hardy Jr, G.T. Reynolds, C.R. Sun, G. Tagliaferri, A.E. Werbrouck and W.H. Moore, Production of Strange Particles by 1.5-Bev Protons in C, Fe, and Pb, Phys. Rev. 119, 2030 (1960),

http://kirkmcd.princeton.edu/examples/EP/bowen_pr_119_2030_60.pdf

26. T. Bowen, J. Hardy Jr, G.T. Reynolds, G. Tagliaferri, A.E. Werbrouck and W.H. Moore, Production of Strange Particles by 2.8-Bev Protons in C, Fe, and Pb, Phys. Rev. 119, 2041 (1960),

http://kirkmcd.princeton.edu/examples/EP/bowen_pr_119_2041_60.pdf

27. G.T. Reynolds, Present Status of Scintillation Chambers, IRE Trans. Nucl. Sci. 7,

115 (1960), http://kirkmcd.princeton.edu/examples/detectors/reynolds_iretns_7_115_60.pdf

28. G.T. Reynolds, R.A. Swanson and D. Scarl, π^{±} Tracks in a Filament Scintillation Chamber, Rev. Sci. Instr. 31, 1011 (1960),

http://kirkmcd.princeton.edu/examples/detectors/reynolds_rsi_31_1011_60.pdf

29. A.E. Werbrouck, Depolarization Due to Target Motion, Phys. Rev. 120, 1463 (1960),

http://kirkmcd.princeton.edu/examples/EP/werbrouck_pr_120_1463_60.pdf

30. A.Z. Hendel, P.A. Piroué, and G.T. Reynolds, Double Cloud-Chamber Investigation of 500 m_{e}, Particles in Cosmic Rays, Phys. Rev. 124, 1580 (1961),

http://kirkmcd.princeton.edu/examples/EP/hendel_pr_124_580_61.pdf

31. S.K. Poultney, G.T. Reynolds, and J.R. Waters, Observation of the Focused Ring of Cerenkov Light from a Single Particle, Rev. Sci. Instr. 33, 574 (1962),

http://kirkmcd.princeton.edu/examples/detectors/poultney_rsi_33_574_62.pdf

32. G.T. Reynolds, J.R. Waters, D.B. Scarl and R.A. Zdanis, A Scintillation Chamber-Image Intensifier Beam Profile Detector, Nucl. Instr. Meth. 17, 44 (1962),

http://kirkmcd.princeton.edu/examples/detectors/waters_nim_17_44_62.pdf

33. J.R. Waters, G.T. Reynolds, D.B. Scarl and R.A. Zdanis, Some Mmeasurements of the Efficiency for Observing Photoelectrons in Image Intensifiers, IRE Trans. Nucl.

Sci. 9, 239 (1962), http://kirkmcd.princeton.edu/examples/detectors/waters_iretns_9_239_62.pdf

34. S.K. Poultney and J. R. Waters, Observation of the Focused Ring of Cerenkov Light from a Single Particle using an Image Intensifier System, IRE Trans. Nucl. Sci. 9, 243

(1962), http://kirkmcd.princeton.edu/examples/detectors/poultney_iretns_9_243_62.pdf

35. G.T. Reynolds, D.B. Scarl, R.A. Swanson, J.R. Waters and R.A. Zdanis, Muon Capture on Carbon, Phys. Rev. 129, 1790 (1963),

http://kirkmcd.princeton.edu/examples/EP/reynolds_pr_129_1790_63.pdf

36. G.T. Reynolds, J.R. Waters and S.K. Poultney, Cherenkov Rings from Single Particles Observed with a High-Gain Image Intensifier System, Nucl. Instr. Meth. 20,

   267 (1963), http://kirkmcd.princeton.edu/examples/detectors/reynolds_nim_20_267_63.pdf

37. After 1963, the Reynolds Group emphasized the physics of bioluminescence.

Fitch Group

1. V. Fitch and R. Motley, Mean Life of K^{+} Mesons, Phys. Rev. 101, 496 (1956),

http://kirkmcd.princeton.edu/examples/EP/fitch_pr_101_496_56.pdf

2. R. Motley and V. Fitch, Mean Life of τ^{+} Mesons, Phys. Rev. 105, 265 (1957),

http://kirkmcd.princeton.edu/examples/EP/fitch_pr_105_265_57.pdf

3. W.K.H. Panofsky, V.L. Fitch, R.M. Motley and W.D. Chesnut, Measurement of the Total Absorption Coefficient of Long-Lived Neutral K Particles, Phys. Rev. 109,

1353 (1958), http://kirkmcd.princeton.edu/examples/EP/panofsky_pr_109_1353_58.pdf

4. V.L. Fitch, Applications of Transistors in Counter Experiments, Proc. HEACC59,

601 (1959), http://kirkmcd.princeton.edu/examples/detectors/fitch_HEACC59_601.pdf

5. V.L Fitch, P.A. Piroué and R.B. Perkins, Mass Difference of Neutral K Mesons,

Nuovo Cim. 22, 1160 (1961), http://kirkmcd.princeton.edu/examples/EP/fitch_nc_22_1160_61.pdf

6. V.L Fitch, S.L. Meyer and P.A. Piroué, Particle Production at Large Angles by 30and 33-Bev Protons Incident on Aluminum and Beryllium, Phys. Rev. 126, 1849

(1962), http://kirkmcd.princeton.edu/examples/EP/fitch_pr_126_1849_62.pdf

7. J.H. Christenson, J.W. Cronin, V.L. Fitch and R. Turlay, K_{1}-K_{2} Mass Difference, Int Conf. Phys. Fund. Asp. Weak Int., 74 (1964),

http://kirkmcd.princeton.edu/examples/EP/christenson_icfawi_74_64.pdf

8. J.H. Christenson, J.W. Cronin, V.L. Fitch and R. Turlay, Studies of coherent regeneration of K_{1}^{0} mesons in various materials, and measurement of the K_{1}^{0}-K_{2}^{0} mass difference, ICHEP64, 107 (1964),

http://kirkmcd.princeton.edu/examples/EP/christenson_ichep64_107.pdf

9. J.H. Christenson, J.W. Cronin, V.L. Fitch and R. Turlay, Evidence for the 2π Decay of the K_{2}^{0} Meson, Phys. Rev. Lett. 13, 138 (1964),

http://kirkmcd.princeton.edu/examples/EP/christenson_prl_13_138_64.pdf

10. V.L. Fitch, R.F. Roth, J.S. Russ, and W. Vernon, Evidence for Constructive Interference Between Coherently Regenerated and CP -Nonconserving Amplitudes, Phys. Rev. Lett. 15, 73 (1965),

http://kirkmcd.princeton.edu/examples/EP/fitch_prl_15_73_65.pdf

11. J.H. Christenson, J.W. Cronin, V.L. Fitch and R. Turlay, Regeneration of K^{0} Mesons and the K_{1}^{0}-K_{2}^{0} Mass Difference, Phys. Rev. 140, B74 (1965),

http://kirkmcd.princeton.edu/examples/EP/christenson_pr_140_B74_65.pdf

12. V.L. Fitch, C.A. Quarles and H.C. Wilkins, Study of the K+ Decay Probability, Phys. Rev. 140, B1088 (1965),

http://kirkmcd.princeton.edu/examples/EP/fitch_pr_140_B1088_65.pdf

13. V.L. Fitch, Experiments in T, C Violation, Proc. ICHEP66, 63 (1966),

    http://kirkmcd.princeton.edu/examples/EP/fitch_ICHEP66_63.pdf

14. V.L. Fitch, R. F. Roth, J. Russ, and W. Vernon, Studies of K_{2}^{0} → π^{+}π^{−} Decay and Interference, Phys. Rev. 164, 1711 (1967),

    http://kirkmcd.princeton.edu/examples/EP/fitch_pr_164_1711_67.pdf

15. V.L. Fitch, C Violation in Eta Decay? Comm. Nucl. Part. Phys. 1(1), 5 (1967),

    http://kirkmcd.princeton.edu/examples/EP/fitch_cnpp_1-1_5_67.pdf

16. V.L. Fitch, Further Results on CP Violations, Comm. Nucl. Part. Phys. 1(2), 47

   (1967), http://kirkmcd.princeton.edu/examples/EP/fitch_cnpp_1-2_47_67.pdf

17. V.L. Fitch, Charge Asymmetries, Comm. Nucl. Part. Phys. 1(4), 117 (1967),

    http://kirkmcd.princeton.edu/examples/EP/fitch_cnpp_1-4_117_67.pdf

18. V.L. Fitch, The Electric Dipole Moment of the Neutron, Comm. Nucl. Part. Phys.

   1(5), 151 (1967), http://kirkmcd.princeton.edu/examples/EP/fitch_cnpp_1-5_151_67.pdf

19. D.F. Bartlett, R.K. Carnegie, V.L. Fitch, K. Goulianos, D.P. Hutchinson, T. Kamae, R.F. Roth, J.S. Russ, and W. Vernon, K_{2}^{0} → 2π^{0} Decay Rate, Phys. Rev. Lett. 21,

   588 (1968), http://kirkmcd.princeton.edu/examples/EP/bartlett_prl_21_558_68.pdf

20. V.L. Fitch, Further Results on CP Violations, Comm. Nucl. Part. Phys. 2(1), 6

   (1968), http://kirkmcd.princeton.edu/examples/EP/fitch_cnpp_2-1_6_68.pdf

21. V.L. Fitch, CP Snow, Comm. Nucl. Part. Phys. 2(3), 63 (1968),

    http://kirkmcd.princeton.edu/examples/EP/fitch_cnpp_2-3_63_68.pdf

22. V.L. Fitch, Weak Interactions-Experimental Progress?, Comm. Nucl. Part. Phys.

   2(supp.), 20 (1968), http://kirkmcd.princeton.edu/examples/EP/fitch_cnpp_2-supp_20_68.pdf

23. R.K. Carnegie, R. Cester, V.L. Fitch, M. Strovink, and L.R. Sulak, K_{1}^{0}-K_{2}^{0} Mass Difference, Phys. Rev. D 4, 1 (1971),

    http://kirkmcd.princeton.edu/examples/EP/carnegie_prd_4_1_71.pdf

24. R.K. Carnegie, R. Cester, V.L. Fitch, M. Strovink, and L.R. Sulak, Self-Contained Determination of the Phase of η_{±}, Phys. Rev. D 6, 2335 (1972),

    http://kirkmcd.princeton.edu/examples/EP/carnegie_prd_6_2335_72.pdf

25. V.L. Fitch, V. Hepp, D. Jensen, M. Strovink, and R.C. Webb, K_{e3}^{0} Charge Asymmetry, Phys. Rev. Lett. 31, 1524 (1973),

    http://kirkmcd.princeton.edu/examples/EP/fitch_prl_31_1524_73.pdf

26. V.L. Fitch, Flower Power and the Energy Doubler/Saver, Fermilab Summer Study,

   87 (1976), http://kirkmcd.princeton.edu/examples/EP/fitch_fnal_summer_study_87_76.pdf

27. R. Cester, V.L. Fitch, R.W. Kadel, R.C. Webb, J.D. Whittaker, M.S. Witherell and M. May, Limits on Production of Charmed Particles by Antiprotons and Pions, Phys.

Rev. Lett. 37, 1178 (1976), http://kirkmcd.princeton.edu/examples/EP/cester_prl_37_1178_76.pdf

28. R. Cester, V.L. Fitch, R.W. Kadel, R.C. Webb, J.D. Whittaker, M.S. Witherell and M. May, Search for Production of D^{∗−} by Pions near Threshold, Phys. Rev. Lett. 40,

139 (1978), http://kirkmcd.princeton.edu/examples/EP/cester_prl_40_139_78.pdf

29. R. Cester, V.L. Fitch, R.W. Kadel, R.C. Webb, J.D. Whittaker and M.S. Witherell, Test of charge-conjugation invariance in pp̄ interactions, Phys. Rev. D 17, 1706

(1978), http://kirkmcd.princeton.edu/examples/EP/cester_prd_17_1706_78.pdf

30. R. Cester, V.L. Fitch M. Isaila, R.W. Kadel and R.C. Webb, Description of a Drift Chamber System Employed in an Experiment at Brookhaven National Laboratory, IEEE Trans. Nucl. Sci. 25, 53 (1978),

http://kirkmcd.princeton.edu/examples/detectors/cester_ieeetns_25_53_78.pdf

31. R. Cester, V.L. Fitch M. Isaila, R.C. Webb and J.D. Whittaker, Design and Performance of a Broadband Focusing Cherenkov Counter, IEEE Trans. Nucl. Sci.

25, 525 (1978), http://kirkmcd.princeton.edu/examples/detectors/cester_ieeetns_25_525_78.pdf

32. V.L. Fitch, A.D. Montag, S.S. Sherman, R.C. Webb, M.S. Witherell, B. Devaux, J. Teiger, R. Turlay, A. Zylberstejn, P. Cavaglia, R. Cester, D. Maurizio, G. Rinaudo, and M. May, Measurement of D^{∗} Production in Pion-Nucleon Interactions at 200 GeV/c, Phys. Rev. Lett. 46, 761 (1981),

http://kirkmcd.princeton.edu/examples/EP/fitch_prl_46_761_81.pdf

33. R. Cester, V.L. Fitch A. Montag, S. Sherman, R.C. Webb, M.S. Witherell, Results on the Performance of a Broad Band Focussing Cherenkov Counter, IEEE Trans. Nucl.

Sci. 28, 425 (1981), http://kirkmcd.princeton.edu/examples/detectors/cester_ieeetns_28_425_81.pdf

34. V.L. Fitch, The discovery of charge-conjugation parity asymmetry, Rev. Mod. Phys.

53, 367 (1981), http://kirkmcd.princeton.edu/examples/EP/fitch_rmp_53_367_81.pdf

35. V.L. Fitch, A.D. Montag, S. Palestini, S. S. Sherman, R.C. Webb, M.S. Witherell, B. Bloch-Devaux, J. Teiger, R. Turlay, A. Zylberstejn, R. Cester, E. Menechetti, G. Rinaudo, D. Maurizio, and M. May, Search for D^{∗} production in pion-nucleon interactions, Phys. Rev. D 33, 1486 (1986),

http://kirkmcd.princeton.edu/examples/EP/fitch_prd_33_1486_86.pdf

36. V.L. Fitch, M.V. Isaila and M.A. Palmer, Limits on the Existence of a Material-Dependent Intermediate-Range Force, Phys. Rev. Lett. 60, 1801 (1988),

http://kirkmcd.princeton.edu/examples/EP/fitch_prl_60_1801_88.pdf

37. V.L. Fitch, The Fifth Force: Fact or Fiction? Blois Conf. (1989),

http://kirkmcd.princeton.edu/examples/EP/fitch_blois_89.pdf

38. J. Belz et al., Search for diffractive dissociation of a long-lived H dibaryon, Phys.

   Rev. D 33, 1486 (1986), http://kirkmcd.princeton.edu/examples/EP/belz_prd_53_R3487_96.pdf

39. V.L. Fitch, Elementary particle physics: The origins, Rev. Mod. Phys. 71, S25

   (1999), http://kirkmcd.princeton.edu/examples/EP/fitch_rmp_71_S25_99.pdf

Cronin Group

Papers by Fitch and Cronin are listed in under the Fitch Group. Papers by Cronin with Princeton authors after he went to U. Chicago are listed under the Piroué Group.

1. J.W. Cronin, Numerical Evaluation of the Pion-Nucleon Forward Scattering Amplitude, Phys. Rev. 118, 824 (1960),

http://kirkmcd.princeton.edu/examples/EP/cronin_pr_118_824_60.pdf

2. B. Cork, L. Kerth, W.A. Wenzel, J.W. Cronin and R.L. Cool, Decay Asymmetry of Σ^{+} and Λ^{0} Hyperons, Phys. Rev. 120, 1000 (1960),

http://kirkmcd.princeton.edu/examples/EP/cork_pr_120_1000_60.pdf

3. J.W. Cronin, Asymmetry in the Pionic Decay Modes of Λ and Σ Hyperons, ICEHP60,

590 (1960), http://kirkmcd.princeton.edu/examples/EP/cronin_ichep60_590.pdf

4. E. Engels, D. Roth, J. Cronin and M. Pyka, Ionization Density Effects in Spark Chambers, IRE Trans. Nucl. Sci. 9, 256 (1962),

http://kirkmcd.princeton.edu/examples/detectors/engels_iretns_9_256_62.pdf

5. J.W. Cronin and O.E. Overseth, Measurement of the Decay Parameters for the Λ^{0} Particle, ICHEP62, 453 (1962),

http://kirkmcd.princeton.edu/examples/EP/cronin_ichep62_453.pdf

6. J.W. Cronin and O.E. Overseth, Measurement of the Decay Parameters of the Λ^{0} Particle, Phys. Rev. 129, 1795 (1963),

http://kirkmcd.princeton.edu/examples/EP/cronin_pr_129_1795_63.pdf

7. A.R. Clark, J.H. Christenson, J.W. Cronin and R. Turlay, Dipion Production at Low Momentum Transfer in π^{−}p Collisions at 1.5 Bev/c, Phys. Rev. 139, B1556 (1965),

http://kirkmcd.princeton.edu/examples/EP/clark_pr_139_B1556_65.pdf

8. J.W. Cronin, P.F. Kunz, W.S. Risk and P.C. Wheeler, Measurement of the Decay Rate of K_{2}^{0} → π^{0} + π^{0}, Phys. Rev. Lett. 18, 25 (1967),

http://kirkmcd.princeton.edu/examples/EP/cronin_prl_18_25_67.pdf

9. P. Basile, J.W. Cronin, B. Thevenet, R. Turlay, S. Zylberajch and A. Zylbersteijn, Study of the Decay Mode K_{L} → πeν, Phys. Lett. B 26, 542 (1968),

http://kirkmcd.princeton.edu/examples/EP/basile_pl_26b_542_68.pdf

10. J.W. Cronin, Weak Interactions and CP Violation–Experimental, ICHEH68, 281

(1968), http://kirkmcd.princeton.edu/examples/EP/cronin_ichep68_281.pdf

11. P. Basile, J.W. Cronin, B. Thevenet, R. Turlay, S. Zylberajch and A. Zylbersteijn, Experimental Study of the Decay K_{L} → π^{+}π^{−}π^{0}, Phys. Lett. B 28, 58 (1968),

http://kirkmcd.princeton.edu/examples/EP/basile_pl_28b_58_68.pdf

12. M. Banner, J.W. Cronin, J.K. Liu, J.E. Pilcher, Measurement of the Branching Ratio (K_{L} → γγ)/(K_{L} → 3π^{0}), Phys. Rev. Lett. 21, 1103 (1968),

    http://kirkmcd.princeton.edu/examples/EP/banner_prl_21_1103_68.pdf

13. M. Banner, J.W. Cronin, J.K. Liu, J.E. Pilcher, Experimental Study of the Decay K_{L} → π^{0}π^{0} and Other Rare Decay Modes, Phys. Rev. 188, 2033 (1969),

    http://kirkmcd.princeton.edu/examples/EP/banner_pr_188_2033_69.pdf

14. P. Basile, J.W. Cronin, B. Thevenet, R. Turlay, S. Zylberajch and A. Zylbersteijn, Experimental Study of the K_{μ3}^{0} Form Factors, Phys. Rev. D 2, 78 (1970),

    http://kirkmcd.princeton.edu/examples/EP/basile_prd_2_78_70.pdf

15. M. Banner, J.W. Cronin, C.M. Hoffman, B.C. Knapp and M.J. Shochet, Measurement of |η_{00}| / |η_{+−}|, Phys. Rev. Lett. 28, 1597 (1972),

    http://kirkmcd.princeton.edu/examples/EP/banner_prl_28_1597_72.pdf

16. M. Banner, J.W. Cronin, C.M. Hoffman, B.C. Knapp and M.J. Shochet, Experimental Study of the Decays K_{S,L} → γγ, Phys. Rev. Lett. 29, 237 (1972),

    http://kirkmcd.princeton.edu/examples/EP/banner_prl_29_237_72.pdf

17. J.W. Cronin, CP symmetry violation—the search for its origin, Rev. Mod. Phys. 53,

   373 (1981), http://kirkmcd.princeton.edu/examples/EP/cronin_rmp_53_373_81.pdf

18. J.W. Cronin and M.S. Greenwood, CP symmetry violation, Phys. Today 35(7), 32

   (1982), http://kirkmcd.princeton.edu/examples/EP/cronin_pt_35-7_38_82.pdf

19. J.W. Cronin, The Experimental Discovery of CP Violation, Lect. Notes Phys 746,

   261 (2008), http://kirkmcd.princeton.edu/examples/EP/cronin_lnp_746_261_08.pdf

20. J.W. Cronin, The discovery of CP violation, Eur. Phys J. H 36, 1 (2011),

    http://kirkmcd.princeton.edu/examples/EP/cronin_epjh_36_487_11.pdf

21. J.W. Cronin, A Life in High-Energy Physics: Success Beyond Expectations, Ann. Rev. Nucl. Part. Sci. 64, 1 (2014),

    http://kirkmcd.princeton.edu/examples/EP/cronin_arnps_64_1_14.pdf

Piroué Group

Includes only papers by the Piroué Group prior to the L3 and CMS colider experiments at CERN. Some early papers by Piroué are listed under the Fitch Group.

1. G. von Dardel, D.H. Frisch, R. Mermod, R.H. Milburn, P.A. Piroué, M. Vivargent, G. Weber and K. Winter, Total Cross Sections for p, p̄, K^{±}, and π^{±} on Hydrogen Between 3 and 10 Gev/c, Phys. Rev. Lett. 5, 333 (1960),

http://kirkmcd.princeton.edu/examples/EP/vondardel_prl_5_333_60.pdf

2. G. von Dardel, R. Mermod, P.A. Piroué, M. Vivargent, G. Weber and K. Winter, Total Cross Sections for Pions on Protons in the Momentum Range 4.5 to 10 Gev/c, Phys. Rev. Lett. 7, 127 (1961),

http://kirkmcd.princeton.edu/examples/EP/vondardel_prl_7_127_61.pdf

3. D.E. Dorfan, J. Eades, L.M. Lederman, W. Lee, C.C. Ting, P. Piroué, S. Smith, J.L. Brown, J.A. Kadyk and G.H. Trilling, Production of Antiprotons via Nuclear Motion, Phys. Rev. Lett. 14, 995 (1965),

http://kirkmcd.princeton.edu/examples/EP/dorfan_prl_14_995_65.pdf

4. P.A. Piroué and A.J.S. Smith, Particle Production by 2.9-BeV Protons Incident on Beryllium and Platinum, Phys. Rev. 148, 1315 (1966),

http://kirkmcd.princeton.edu/examples/EP/piroue_pr_148_1315_66.pdf

5. W.T. Ford, A. Lemonick, U. Nauenberg and P.A. Piroué, Comparison of the K^{+} and K^{−} Decay Rates into the τ and K_{μ2} Modes, Phys. Rev. Lett. 18, 1214 (1967),

http://kirkmcd.princeton.edu/examples/EP/ford_prl_18_1214_67.pdf

6. W.J. Hogan, P.A. Piroué and A.J.S. Smith, K^{+}-Meson Production in p-p Collisions at 2.5-3.0 GeV, Phys. Rev. 166, 1472 (1968),

http://kirkmcd.princeton.edu/examples/EP/hogan_pr_166_1472_68.pdf

7. W.T. Ford, P.A. Piroué, R.S. Remmel, A.J.S. Smith and P.A. Souder, Wire-spark-chamber system with on-line computers, Nucl. Instr. Meth. 87, 241

(1970), http://kirkmcd.princeton.edu/examples/detectors/ford_nim_87_241_70.pdf

8. W.T. Ford, P.A. Piroué, R.S. Remmel, A.J.S. Smith and P.A. Souder, Search for Violation of CP Invariance in τ^{±} Decay, Phys. Rev. Lett. 25, 1370 (1970),

http://kirkmcd.princeton.edu/examples/EP/ford_prl_25_1370_70.pdf

9. M. Bourquin, J.P. Boymond, P. Extermann, J. Marasco, R. Mermod, P.A. Piroué, H. Suter, P Basile, S. Brehin, A. Diamant-Berger, P. Kunz, M. Lemoine, R. Turlay and A. Zylbersztejn, Measurement of K_{e4}^{+}-decay rates and test of the ΔS = ΔQ rule, Phys. Lett. B 36, 615 (1971),

http://kirkmcd.princeton.edu/examples/EP/bourquin_pl_36b_615_71.pdf

10. P Basile, S. Brehin, A. Diamant-Berger, P. Kunz, M. Lemoine, R. Turlay, A. Zylbersztejn, M. Bourquin, J.P. Boymond, P. Extermann, J. Marasco, R. Mermod, P.A. Piroué and H. Suter, Determination of the low-energy π-π phase shifts and form factors in K_{e4}^{+} decay, Phys. Lett. B 36, 619 (1971),

http://kirkmcd.princeton.edu/examples/EP/basile_pl_36b_619_71.pdf

11. W.T. Ford, P.A. Piroué, R.S. Remmel, A.J.S. Smith and P.A. Souder, Experimental study on the τ^{±} decay matrix element, Phys. Lett. B 38, 335 (1972),

http://kirkmcd.princeton.edu/examples/EP/ford_pl_38b_335_72.pdf

12. J.W. Cronin, H.J. Frisch, M.J. Shochet, J.P. Boymond, P.A. Piroué and R.L. Sumner, Production of Hadrons with Large Transverse Momentnm at 200 and 300 GeV, Phys. Rev. Lett. 31, 1426 (1973),

http://kirkmcd.princeton.edu/examples/EP/cronin_prl_31_1426_73.pdf

13. J.P. Boymond, R. Mermod, P.A. Piroué, R.L. Sumner,J.W. Cronin, H.J. Frisch, M.J. Shochet, Observation of Large-Transverse-Momentum Muons Directly Produced by 300-Gev Protons, Phys. Rev. Lett. 33, 112 (1974),

http://kirkmcd.princeton.edu/examples/EP/boymond_prl_33_112_74.pdf

14. J.W. Cronin, H.J. Frisch, M.J. Shochet, J.P. Boymond, R. Mermod, P.A. Piroué and R.L. Sumner, Search for Massive Penetrating Particles Produced by 300-GeV Protons, Phys. Rev. D 10, 3093 (1974),

http://kirkmcd.princeton.edu/examples/EP/cronin_prd_10_3093_74.pdf

15. J.W. Cronin, H.J. Frisch, M.J. Shochet, J.P. Boymond, R. Mermod, P.A. Piroué and R.L. Sumner, Atomic Number Dependence of Hadron Production at Larse Transverse Momentum in 300 GeV Proton-Nucleus Collisions, ICHEP74 (1974),

http://kirkmcd.princeton.edu/examples/EP/cronin_fermilab-conf-74-147.pdf

16. P.A. Piroué, Leptons at High P_{T}, ICHEP74, V47 (1974),

http://kirkmcd.princeton.edu/examples/EP/piroue_ichep74_v47.pdf

17. J.W. Cronin, H.J. Frisch, M.J. Shochet, J.P. Boymond, P.A. Piroué and R.L. Sumner, Production of hadrons with large transverse momentum at 200, 300, and 400 GeV, Phys. Rev. D 11, 3105 (1975),

http://kirkmcd.princeton.edu/examples/EP/cronin_prd_11_3105_75.pdf

18. L. Kluberg, P.A. Piroué, R.L. Sumner, D. Antreaysan, J.W. Cronin, H.J. Frisch and M.J. Shochet, Production of Massive Muon Pairs by 300- and 400-Gev Protons, Phys. Rev. Lett. 37, 1451 (1976),

http://kirkmcd.princeton.edu/examples/EP/kluberg_prl_37_1451_76.pdf

19. D. Antreaysan, J.W. Cronin, H.J. Frisch, M.J. Shochet, L. Kluberg, P.A. Piroué and R.L. Sumner, Production of π^{+}+ and π^{−} at Large Transverse Momentum in p-p and p-d Collisions at 200, 300, and 400 GeV, Phys. Rev. Lett. 38, 112 (1977),

http://kirkmcd.princeton.edu/examples/EP/antreasyan_prl_38_112_77.pdf

20. D. Antreaysan, J.W. Cronin, H.J. Frisch, M.J. Shochet, L. Kluberg, P.A. Piroué and R.L. Sumner, Production of Kaoans, Protons, and Antiproton at Large Transverse Momentum in p-p and p-d Collisions at 200, 300, and 400 GeV, Phys. Rev. Lett. 38,

115 (1977), http://kirkmcd.princeton.edu/examples/EP/antreasyan_prl_38_115_77.pdf

21. L. Kluberg, P.A. Piroué, R.L. Sumner, D. Antreaysan, J.W. Cronin, H.J. Frisch and M.J. Shochet, Atomic-Number Dependence of Large-Transverse-Momentum Hadron Production by Protons, Phys. Rev. Lett. 38, 670 (1977),

http://kirkmcd.princeton.edu/examples/EP/kluberg_prl_38_670_77.pdf

22. D. Antreaysan, J.W. Cronin, H.J. Frisch, M.J. Shochet, L. Kluberg, J.J. Mueller, P.A. Piroué and R.L. Sumner, Test of Scaling in Muon-Pair Production by Hadrons, Phys. Rev. Lett. 39, 906 (1977),

http://kirkmcd.princeton.edu/examples/EP/antreaysan_prl_39_906_77.pdf

23. D. Antreaysan, G. Coconni, J.W. Cronin, H.J. Frisch, K. Olive, M.J. Shochet, L. Kluberg, J. Mueller, P.A. Piroué and R.L. Sumner, Search for Quarks Produced with Large Transverse Momentum in 400-GeV Proton-Nucleus Collisions, Phys. Rev.

Lett. 39, 513 (1977), http://kirkmcd.princeton.edu/examples/EP/antreasyan_prl_39_513_77.pdf

24. D. Antreaysan, J.W. Cronin, H.J. Frisch, M.J. Shochet, L. Kluberg, P.A. Piroué and R.L. Sumner, Production of hadrons at large transverse momentum in 200-, 300-, and 400-GeV p-p and p-nucleus collisions, Phys. Rev. D 19, 764 (1979),

http://kirkmcd.princeton.edu/examples/EP/antreaysan_prd_19_764_79.pdf

25. H.J. Frisch, N.D. Giokaris, J.M. Green, C. Grosso-Pilcher, G. Hanson, K.F. Johnson, M.D. Mestayer, L. Schachinger, M.J. Shochet, M.L. Swartz, A.M. Hailing, P.A. Piroué, B.G. Pope, and R.L. Sumner, Relative Production of π^{+}, K^{+}, p, and p̄ at Large Transverse Momentum in 200- and 300-GeV π^{−}-p Collisions, Phys. Rev. Lett.

44, 511 (1980), http://kirkmcd.princeton.edu/examples/EP/frisch_prl_44_511_80.pdf

26. N.D. Giokaris, J.M. Green, H.B. Greenlee, C. Grosso-Pilcher, M.D. Mestayer, L. Schachinger, M.J. Shochet, M.L. Swartz, P.A. Piroué, B.G. Pope and R.L. Sumner, Comparison of the Production of Charged Hadrons at Large Transverse Momentum in π^{−}p and pp Collisions, Phys. Rev. Lett. 47, 1690 (1981),

http://kirkmcd.princeton.edu/examples/EP/giokaris_prl_47_1690_81.pdf

27. H.J. Frisch, N.D. Giokaris, J.M. Green, C. Grosso-Pilcher, M.D. Mestayer, L. Schachinger, M.J. Shochet, M.L. Swartz, P.A. Piroué, B.G. Pope, R.M. Rohm, D.P. Stickland, R.L. Sumner and C. Whitmer, Atomic-weight dependence of muon-pair production in 225-GeV/c π^{−}-nucleus interactions, Phys. Rev. D 25, 2000

(1982), http://kirkmcd.princeton.edu/examples/EP/frisch_prd_25_2000_82.pdf

28. H.J. Frisch, N.D. Giokaris, J.M. Green, C. Grosso-Pilcher, M.D. Mestayer, L. Schachinger, M.J. Shochet, P.A. Piroué, B.G. Pope and R.L. Sumner, Inclusive production of hadrons at high P_{T} in 200 and 300 GeV π^{−}p and π^{−}-nucleus collisions, Phys. Rev. D 27, 1001 (1983),

http://kirkmcd.princeton.edu/examples/EP/frisch_prd_27_1001_83.pdf

29. M.L. Swartz, H.J. Frisch, H.B. Greenlee, C. Grosso-Pilcher, K.F. Johnson, M.D. Mestayer, L. Schachinger, M.J. Shochet, P.A. Piroué, B.G. Pope, D.P. Stickland and R.L. Sumner, Atomic-Mass Dependence of the Transverse-Momentum Distribution of Massive Muon Pairs from 225-Gev/c π^{−}-Nucleus Collisions, Phys. Rev. Lett. 53, 32 (1984),

http://kirkmcd.princeton.edu/examples/EP/swartz_prl_53_32_84.pdf

30. H.B. Greenlee, H.J. Frisch, C. Grosso-Pilcher, K.F. Johnson, M.D. Mestayer, L. Schachinger, M.J. Shochet, M.L. Swartz, P.A. Piroué, B.G. Pope, D.P. Stickland and R.L. Sumner, Production of Massive Muon Pairs in π^{−}-Nucleus Collisions, Phys.

Rev. Lett. 55, 1555 (1985), http://kirkmcd.princeton.edu/examples/EP/greenlee_prl_55_1555_85.pdf

Smith Group

Early papers by Smith at Princeton are listed in under the Piroué Group. The list below is of papers prior to the BaBar experiment at SLAC.

1. K.J. Anderson, G.G. Henry, K.T. McDonald, J.E. Pilcher, E.I. Rosenberg, J.G. Branson, G.H. Sanders, A.J.S. Smith and J.J. Thaler, Production of Muon Pairs by 150-GeV/c π^{+} and Protons, Phys. Rev. Lett. 36, 237 (1976),

http://kirkmcd.princeton.edu/examples/EP/anderson_prl_36_237_76.pdf

2. J.G. Branson, G.H. Sanders, A.J.S. Smith, J.J. Thaler, K.J. Anderson, G.G. Henry, K.T. McDonald, J.E. Pilcher and E.I. Rosenberg, Dimuon Production by 150 Gev/c π^{+} and Protons in a Large Acceptance Detector, DPF Conf. (Seattle, August 1975),

http://kirkmcd.princeton.edu/examples/EP/branson_fermilab-conf-75-193.pdf

3. K.J. Anderson, G.G. Henry, K.T. McDonald, C. Newamn, J.E. Pilcher, E.I. Rosenberg, J.G. Branson, G.E. Hogan, R.D. Pisarski, G.H. Sanders, A.J.S. Smith and J.J. Thaler, Production of the J(3.1) and ψ′(3.7) by 225 GeV π^{+}, π^{−} and

Protons, ICHEP76, http://kirkmcd.princeton.edu/examples/EP/anderson_fermilab-conf-76-166.pdf

4. K.J. Anderson, G.G. Henry, K.T. McDonald, J.E. Pilcher, E.I. Rosenberg, J.G. Branson, G.H. Sanders, A.J.S. Smith and J.J. Thaler, Inclusive μ-Pair Production at 150 GeV by π^{+} Mesons and Protons, ICHEP76, B162,

kirkmcd.princeton.edu/papers/anderson_fermilab-conf-76-167-e.pdf

5. K.J. Anderson, G.G. Henry, K.T. McDonald, C. Newamn, J.E. Pilcher, E.I. Rosenberg, J.G. Branson, G.E. Hogan, R.D. Pisarski, G.H. Sanders, A.J.S. Smith and J.J. Thaler, Production of Continuum Muon Pairs at 225 GeV by Pions and

Protons, ICHEP76, http://kirkmcd.princeton.edu/examples/EP/anderson_fermilab-conf-76-170.pdf

6. K.J. Anderson, G.G. Henry, K.T. McDonald, J.E. Pilcher, E.I. Rosenberg, J.G. Branson, G.H. Sanders, A.J.S. Smith and J.J. Thaler, Contribution of Muon Pairs to the Yield of Single Prompt Muons, Phys. Rev. Lett. 37, 803 (1976),

http://kirkmcd.princeton.edu/examples/EP/anderson_prl_37_803_76.pdf

7. J.G. Branson, G.H. Sanders, A.J.S. Smith, J.J. Thaler, K.J. Anderson, G.G. Henry, K.T. McDonald, J.E. Pilcher and E.I. Rosenberg, Observation of Prompt Single Muons and Dimuons in Hadron-Nucleus Collisions at 200 GeV/c, Phys. Rev. Lett.

38, 457 (1977), http://kirkmcd.princeton.edu/examples/EP/branson_prl_38_457_77.pdf

8. J.G. Branson, G.H. Sanders, A.J.S. Smith, J.J. Thaler, K.J. Anderson, G.G. Henry, K.T. McDonald, J.E. Pilcher and E.I. Rosenberg, Search for Muons Produced in Conjunction with the J/ψ Particle, Phys. Rev. Lett. 38, 580 (1977),

http://kirkmcd.princeton.edu/examples/EP/branson_prl_38_580_77.pdf

9. J.G. Branson, G.H. Sanders, A.J.S. Smith, J.J. Thaler, K.J. Anderson, G.G. Henry, K.T. McDonald, J.E. Pilcher and E.I. Rosenberg, Production of the J/ψ and ψ′(3.7) by 225-GeV/c π^{±} and Proton Beams on C and Sn Targets, Phys. Rev. Lett. 38, 1331

(1977), http://kirkmcd.princeton.edu/examples/EP/branson_prl_38_1331_77.pdf

10. J.G. Branson, G.H. Sanders, A.J.S. Smith, J.J. Thaler, K.J. Anderson, G.G. Henry, K.T. McDonald, J.E. Pilcher and E.I. Rosenberg, Hadronic Production of Massive Muon Pairs: Dependence on Incident-Particle Type and on Target Nucleus, Phys. Rev. Lett. 38, 1334 (1977),

http://kirkmcd.princeton.edu/examples/EP/branson_prl_38_1334_77.pdf

11. K.J. Anderson, R.N. Coleman, G.E. Hogan, K.P. Karhi, K.T. McDonald, C.B. Newman, J.E. Pilcher, E.I. Rosenberg, G.H. Sanders, A.J.S. Smith and J.J. Thaler, Production of Muon Pairs by 225-GeVc π^{±}, K^{±}, p^{±} Beams on Nuclear Targets, Phys. Rev. Lett. 42, 944 (1979),

http://kirkmcd.princeton.edu/examples/EP/anderson_prl_42_944_79.pdf

12. G.E. Hogan, K.J. Anderson, R.N. Coleman, K.P. Karhi, K.T. McDonald, C.B. Newman, J.E. Pilcher, E.I. Rosenberg, G.H. Sanders, A.J.S. Smith and J.J. Thaler, Comparison of Muon-Pair Production to the Quark-Antiquark Annihilation Model, Phys. Rev. Lett. 42, 948 (1979),

http://kirkmcd.princeton.edu/examples/EP/hogan_prl_42_948_79.pdf

13. C.B. Newman, K.J. Anderson, R.N. Coleman, G.E. Hogan, K.P. Karhi, K.T. McDonald, J.E. Pilcher, E.I. Rosenberg, G.H. Sanders, A.J.S. Smith and J.J. Thaler, Determination of the Pion Structure Function from Muon-Pair Production, Phys. Rev. Lett. 42, 951 (1979),

http://kirkmcd.princeton.edu/examples/EP/newman_prl_42_951_79.pdf

14. K.J. Anderson, R.N. Coleman, K.P. Karhi, C.B. Newman, J.E. Pilcher, E.I. Rosenberg, G.E. Hogan, K.T. McDonald, G.H. Sanders, A.J.S. Smith and J.J. Thaler, Evidence for Longitudinal Photon Polarization in Muon-Pair Production by Pions, Phys. Rev. Lett. 42, 951 (1979),

http://kirkmcd.princeton.edu/examples/EP/anderson_prl_43_1219_79.pdf

15. K.J. Anderson, R.N. Coleman, K.P. Karhi, C.B. Newman, J.E. Pilcher, E.I. Rosenberg, J.J. Thaler, G.E. Hogan, K.T. McDonald, G.H. Sanders and A.J.S. Smith, Search for additional muons in hadronic production of J/ψ particles, Phys. Rev. D 21, 3075 (1980),

http://kirkmcd.princeton.edu/examples/EP/anderson_prd_21_3075_90.pdf

16. R.N. Coleman, K.J. Anderson, K.P. Karhi, C.B. Newman, J.E. Pilcher, E.I. Rosenberg, J.J. Thaler, G.E. Hogan, K.T. McDonald, G.H. Sanders and A.J.S. Smith, Search for additional muons in hadronic production of J/ψ particles, Phys. Rev. Lett. 44, 1313 (1980),

http://kirkmcd.princeton.edu/examples/EP/coleman_prl_44_1313_80.pdf

17. I-H. Chiang, R.A. Johnson, B. Kwan, T.F. Kycia, K.K. Li, L.S. Littenberg, A. Wijangco, L.A. Garren, J.J. Thaler, G.E. Hogan, K.T. McDonald and A.J.S. Smith, A Search for Narrow States Produced in the Reation π^{−}p → nγ’s at 13-GeV/c, AIP Conf. Proc. 67, 415 (1981),

http://kirkmcd.princeton.edu/examples/EP/chiang_aipcp_67_415_81.pdf

18. C.E. Adolphsen, J.P. Alexander, K.J. Anderson, J. Conway, J. Heinrich, K.W. Merrit, J.E. Pilcher, J. Elias, E.I. Rosenberg, D.T. Simpson, C. Biino, J.F. Greenhalgh, W.C. Louis, K.T. McDonald, S. Palestini, F.C. Shoemaker and A.J.S. Smith, Status of E-615 at Fermilab: Production of Muon Pairs in the Forward Direction, Proc. Drell-Yan Workshop (Fermilab, 1982), p. 271,

http://kirkmcd.princeton.edu/examples/EP/adophsen_fnal-dyw_271_82.pdf

19. I-H. Chiang, R.A. Johnson, B. Kwan, T.F. Kycia, K.K. Li, L.S. Littenberg, A. Wijangco, A.M. Halling, G.E. Hogan, K.T. McDonald, A.J.S. Smith, M.H. Ye, L.A. Garren and J.J. Thaler, Experimental Search for Narrow Resonances in the Reaction π^{−}p → γγn at 13 GeV/c, Phys. Lett. B 140, 145 (1984),

http://kirkmcd.princeton.edu/examples/EP/chiang_pl_140b_145_84.pdf

20. C.E. Adolphsen, J.P. Alexander, K.J. Anderson, J. Conway, J. Heinrich, K.W. Merrit, J.E. Pilcher, E.I. Rosenberg, D.T. Simpson, C. Biino, J.F. Greenhalgh, W.C. Louis, K.T. McDonald, S. Palestini, F.C. Shoemaker and A.J.S. Smith, Measurement of muon pair production near x_{F} = 1, ICHEP84, 271 (Leipzig, 1983),

http://kirkmcd.princeton.edu/examples/EP/anderson_ichep84_279.pdf

21. S. Palestini, C. Biino, J.F. Greenhalgh, W.C. Louis, K.T. McDonald, F.C. Shoemaker, A.J.S. Smith, C.E. Adolphsen, J.P. Alexander, K.J. Anderson, J. Conway, J.G. Heinrich, K.W. Merrit, J.E. Pilcher, E.I. Rosenberg and D.T. Simpson, Pion Structure as Observed in the Reaction π^{−}N → μ^{+}μ^{−}X at 80 GeV/c, Phys. Rev. Lett. 55, 2649 (1985),

http://kirkmcd.princeton.edu/examples/EP/palestini_prl_55_2649_85.pdf

22. C. Biino, J.F. Greenhalgh, W.C. Louis, K.T. McDonald, S. Palestini, F.C. Shoemaker, A.J.S. Smith, C.E. Adolphsen, J.P. Alexander, K.J. Anderson, J. Conway, J.G. Heinrich, K.W. Merrit, J.E. Pilcher, E.I. Rosenberg and D.T. Simpson, An Apparatus to Measure the Structure of the Pion, Nucl. Instr. Meth. A 243, 323 (1985),

http://kirkmcd.princeton.edu/examples/detectors/biino_nim_a243_323_86.pdf

23. W.C. Louis, C.E. Adolphsen, J.P. Alexander, K.J. Anderson, C. Biino, J. Conway, J.F. Greenhalgh, J.G. Heinrich, K.T. McDonald, K.W. Merrit, J.E. Pilcher, E.I. Rosenberg, S. Palestini, F.C. Shoemaker, D.T. Simpson and A.J.S. Smith, Upper Limits on the Decay D^{0} → μ + μ− and on D^{0}-D̄^{0} Mixing, Phys. Rev. Lett. 56, 1027

(1986), http://kirkmcd.princeton.edu/examples/EP/louis_prl_56_1027_86.pdf

24. J.P. Alexander, C.E. Adolphsen, K.J. Anderson, J. Conway, J. Heinrich, K.W. Merrit, J.E. Pilcher, E.I. Rosenberg, D.T. Simpson, C. Biino, J.F. Greenhalgh, W.C. Louis, K.T. McDonald, S. Palestini, F.C. Shoemaker and A.J.S. Smith, Longitudinal photon polarization in muon pair production at high x_{F}, Phys. Rev. D

34, 315 (1986), http://kirkmcd.princeton.edu/examples/EP/alexander_prd_34_315_86.pdf

25. I-H. Chiang, R.A. Johnson, B. Kwan, T.F. Kycia, K.K. Li, L.S. Littenberg, A. Wijangco, A.M. Halling, G.E. Hogan, J.c. Linini, C.G. Lu, K.T. McDonald, A.J.S. Smith, M.H. Ye, L.A. Garren and J.J. Thaler, Search for exclusive J/ψ production, Phys. Rev. D 34, 1619 (1986),

http://kirkmcd.princeton.edu/examples/EP/chiang_prd_34_1619_86.pdf

26. C. Biino, J.F. Greenhalgh, P. Kaaret, W.C. Louis, K.T. McDonald, S. Palestini, F.C. Shoemaker, A.J.S. Smith, C.E. Adolphsen, J.P. Alexander, K.J. Anderson, J. Conway, J.G. Heinrich, J.E. Pilcher, A. Possoz, and E.I. Rosenberg, J/ψ Longitudinal Polarization from πN Interactions, Phys. Rev. Lett. 58, 2523 (1987),

http://kirkmcd.princeton.edu/examples/EP/biino_prl_58_2523_87.pdf

27. J.S. Conway, C.E. Adolphsen, J.P. Alexander, K.J. Anderson, J.G. Heinrich, J.E. Pilcher, A. Possoz, E.I. Rosenberg, C. Biino, J.F. Greenhalgh, W.C. Louis, K.T. McDonald, S. Palestini, F.C. Shoemaker and A.J.S. Smith, Experimental study of muon pairs produced by 252-GeV pions on tungsten, Phys. Rev. D 39, 92 (1989),

http://kirkmcd.princeton.edu/examples/EP/conway_prd_39_92_89.pdf

28. J.G. Heinrich, C.E. Adolphsen, J.P. Alexander, K.J. Anderson, J.S. Conway, J.E. Pilcher, A. Possoz, E.I. Rosenberg, C. Biino, J.F. Greenhalgh, W.C. Louis, K.T. McDonald, S. Palestini, F.C. Shoemaker and A.J.S. Smith, Measurement of the Ratio of Sea to Valence Quarks in the Nucleon, Phys. Rev. Lett. 63, 356 (1989),

http://kirkmcd.princeton.edu/examples/EP/heinrich_prl_63_356_89.pdf

29. M.S. Atiya et al., A Study of K^{+} → π^{+}νν̄, Nucl. Phys. B (Proc. Suppl.) 8, 364

(1989), http://kirkmcd.princeton.edu/examples/EP/atiya_npbps_8_364_89.pdf

30. M.S. Atiya et al., Search for a Light Higgs Boson in the Decay K^{+} → π^{+}H, H → μ + μ−, Phys. Rev. Lett. 63, 2177 (1989),

http://kirkmcd.princeton.edu/examples/EP/atiya_prl_63_2177_89.pdf

31. K.T. McDonald, Pion and Nucleon Structure as Probed in the Reaction π^{±}N → μ^{+}μ^{−}X at 253 GeV, Nucl. Phys. B (Proc. Suppl.) 7B, 104 (1989),

kirkmcd.princeton.edu/papers/mcdonald_npb(ps)_7b_104_89.pdf

32. M.S. Atiya et al., Search for the Decay K^{+} → π^{+}νν̄, Phys. Rev. Lett. 64, 21 (1990),

http://kirkmcd.princeton.edu/examples/EP/atiya_prl_64_21_90.pdf

33. M.S. Atiya et al., Search for the Decay K^{+} → π^{+}γγ, Phys. Rev. Lett. 65, 1188

(1990), http://kirkmcd.princeton.edu/examples/EP/atiya_prl_65_1188_90.pdf

34. M.S. Atiya et al., Upper Limit on the Branching Ratio for the Decay π^{0} → νν̄, Phys.

Rev. Lett. 66, 2189 (1991), http://kirkmcd.princeton.edu/examples/EP/atiya_prl_66_2189_91.pdf

35. M.S. Atiya et al., Recent results on rare K^{+} decays from BNL E787, AIP Conf. Proc.

243, 615 (1992), http://kirkmcd.princeton.edu/examples/EP/atiya_aipcp_243_615_92.pdf

36. M.S. Atiya et al., A Search for K^{+} → π^{+}νν̄ and Other Rare K^{+} Decays, Nucl. Phys.

A 527, 727c (1991), http://kirkmcd.princeton.edu/examples/EP/atiya_np_a527_727c_91.pdf

37. M.S. Atiya et al., Search for the Decay π^{0} → γ + X, Phys. Rev. Lett. 69, 733 (1992),

http://kirkmcd.princeton.edu/examples/EP/atiya_prl_69_733_92.pdf

38. M.S. Atiya et al., A detector to search for K^{+} → π^{+}νν̄, Nucl. Instr. Meth. A 321, 192

(1992), http://kirkmcd.princeton.edu/examples/detectors/atiya_nim_a321_192_92.pdf

39. M.S. Atiya et al., Search for the Decay K^{+} → π^{+}νν̄, Phys. Rev. Lett. 70, 2521

(1993), http://kirkmcd.princeton.edu/examples/EP/atiya_prl_70_2521_93.pdf

40. M.S. Atiya et al., Search for the decays K^{+} → π^{+}νν̄ and K^{+} → π^{+}X^{0} for 150 < M_{X^{0}} < 250 MeV/c^{2}, Phys. Rev. D 48, R1 (1993),

http://kirkmcd.princeton.edu/examples/EP/atiya_prd_48_R1_93.pdf

41. J.G. Heinrich, C. Biino, J.F. Greenhalgh, W.C. Louis, K.T. McDonald, S. Palestini, D.P. Russell, F.C. Shoemaker and A.J.S. Smith C.E. Adolphsen, J.P. Alexander, K.J. Anderson, J.S. Conway, J.E. Pilcher, A. Possoz, and E.I. Rosenberg, Higher Twist Effects in the Reaction π^{−}N → μ^{+}μ^{−}X at 253 GeV/c, Phys. Rev. D 44, 1909

(1991), http://kirkmcd.princeton.edu/examples/EP/heinrich_prd_44_1909_91.pdf

42. S. Adler et al., Search for the Decay K^{+} → π^{+}νν̄, Phys. Rev. Lett. 76, 1421 (1996),

http://kirkmcd.princeton.edu/examples/EP/adler_prl_76_1421_96.pdf

43. P. Kitching et al., Observation of the Decay K^{+} → π^{+}γγ, Phys. Rev. Lett. 79, 4079

(1997), http://kirkmcd.princeton.edu/examples/EP/kitching_prl_79_4079_97.pdf

44. S. Adler et al., Evidence for the Decay K^{+} → π^{+}νν̄, Phys. Rev. Lett. 79, 2204

(1997), http://kirkmcd.princeton.edu/examples/EP/adler_prl_79_2204_97.pdf

45. S. Adler et al., Upper limit on the decay K^{+} → e^{+}νμ^{+}μ^{−}, Phys. Rev. D 58, 012003

(1998), http://kirkmcd.princeton.edu/examples/EP/adler_prd_58_012003_98.pdf

46. S. Adler et al., Further Search for the Decay K^{+} → π^{+}νν̄, Phys. Rev. Lett. 84, 3768

(2000), http://kirkmcd.princeton.edu/examples/EP/adler_prl_84_3768_00.pdf

47. S. Adler et al., Measurement of Structure-Dependent K^{+} → μ^{+}ν_{μ}γ Decay, Phys.

Rev. Lett. 85, 2256 (2000), http://kirkmcd.princeton.edu/examples/EP/adler_prl_85_2256_00.pdf

48. S. Adler et al., Measurement of Direct Photon Emission in K^{+} → π^{+}π^{0}γ Decay, Phys. Rev. Lett. 85, 4586 (2000),

http://kirkmcd.princeton.edu/examples/EP/adler_prl_85_4586_00.pdf

49. S. Adler et al., Search for the decay K^{+} → π^{+}π^{0}νν̄, Phys. Rev. D 63, 032004 (2001),

    http://kirkmcd.princeton.edu/examples/EP/adler_prd_63_032004_01.pdf

50. S. Adler et al., Search for the rare decay K^{+} → π^{+}γ, Phys. Rev. D 65, 052009

   (2002), http://kirkmcd.princeton.edu/examples/EP/adler_prd_65_052009_02.pdf

51. S. Adler et al., Further Evidence for the Decay K^{+} → π^{+}νν̄, Phys. Rev. Lett. 88,

   041803 (2002), http://kirkmcd.princeton.edu/examples/EP/adler_prl_88_041803_02.pdf

52. S. Adler et al., Search for the decay K^{+} → π^{+}νν̄ in the momentum region P_{π} < 195 MeV/c, Phys. Lett. B 537, 211 (2002),

    http://kirkmcd.princeton.edu/examples/EP/adler_pl_b537_211_02.pdf

53. S. Adler et al., Further search for the decay K^{+} → π^{+}νν̄ in the momentum region P < 195 MeV/c, Phys. Rev. D 70, 037102 (2004),

    http://kirkmcd.princeton.edu/examples/EP/adler_prd_70_037102_04.pdf

54. S. Adler et al., Measurement of the K^{+} → π^{+}νν̄ branching ratio, Phys. Rev. D 77,

   052003 (2008), http://kirkmcd.princeton.edu/examples/EP/adler_prd_77_052003_08.pdf

55. S. Adler et al., Measurement of the K^{+} → π^{0}μ^{+}ν_{μ}γ branching ratio, Phys. Rev. D 81,

   092001 (2010), http://kirkmcd.princeton.edu/examples/EP/adler_prd_81_092001_10.pdf

McDonald Group – Detector Physics

Papers by the McDonald Group fall into several catergories, which will be listed in separate sections of this bibliography.

1. Studies of the Drell-Yan process and pion structure; these appear under the Smith Group.

2. Dectector Physics, including efforts towards the Bottom Collider Detector at the SSC.

3. Accelerator Physics, other than efforts towards a Muon Collider and Neutrino Factory.

4. BaBar experiment at SLAC; not listed here.

4. Muon Collider and Neutrino Factory.

5. DayaBay Reactor Neutrino Experiment.

McDonald’s complete bibliography is at http://kirkmcd.princeton.edu/tex/bib.pdf

1. K.T. McDonald and J.J Thaler, A Lepton Detector Facility for ISABELLE, BNL ISABELLE Workshop (1976),

http://kirkmcd.princeton.edu/examples/detectors/mcdonald_isabelle_77.pdf

2. G.H. Sanders, S. Sherman, K.T. McDonald, A.J.S. Smith and J.J Thaler, Drift Chamber Performance in the Field of a Superconducting Magnet: Measurement of the Drift Angle, IEEE Trans. Nucl. Sci. 25, 56 (1978),

http://kirkmcd.princeton.edu/examples/detectors/sanders_ieeetns_25_56_78.pdf

3. G.H. Sanders, S. Sherman, K.T. McDonald, A.J.S. Smith and J.J Thaler, Drift Chamber Performance in a Strong Magnetic Field: Measurement of the Drift Angle up to 4.5 T, Nucl. Instr. and Meth. 156, 159 (1978),

http://kirkmcd.princeton.edu/examples/detectors/sanders_nim_156_159_78.pdf

4. K.T. McDonald and A.H. Walenta, Performance of the Parallel Multiplate Avalanche Chamber (in Chinese), Physica Energie Fortis et Physics Nuclearis 7, 285 (1983),

http://kirkmcd.princeton.edu/examples/detectors/mcdonald_pefpn_7_285_83.pdf

5. K.T. McDonald, Prospects for Beauty Physics at Hadron Colliders, A.I.P. Conf.

Proc. 185, 526 (1989), kirkmcd.princeton.edu/examples/detectors/mcdonald_aipcp_185_526_89.pdf

6. K.T. McDonald, Prospects for Beauty Physics at Hadron Colliders, Ann. N.Y. Acad.

Sci. 578, 215 (1989), http://kirkmcd.princeton.edu/examples/detectors/mcdonald_apny_578_215_89.pdf

7. H. Castro et al.(presented by K.T. McDonald), The Bottom Collider Detector, in Physics at Fermilab in the 1990’s, D. Green and H. Lubatti, eds., (Breckenridge,

Aug. 1989), p. 287, kirkmcd.princeton.edu/bphys/breck_0889.pdf

8. W.S. Anderson, J.C. Armitage, E. Dunn, J.G. Heinrich, C. Lu, K.T. McDonald, J. Weckel and Y. Zhu, Electron Attachment, Effective Ionization Coefficient, and Electron Drift Velocity for CF_{4} Gas Mixtures, Nucl. Instr. and Meth. A 323, 273

(1992), http://kirkmcd.princeton.edu/examples/accel/anderson_nim_a323_273_92.pdf

9. N.S. Lockyer, J.E. Millan, C. Lu, K.T. McDonald and A. Lopez, Observation of Čerenkov Rings Using a Low-Pressure Parallel-Plate Chamber and a Solid Cesium-Iodide Photocathode, Nucl. Instr. and Meth. A 332, 142 (1993),

http://kirkmcd.princeton.edu/examples/accel/lockyer_nim_a332_142_93.pdf

10. C. Lu, K.T. McDonald and Y. Zhu, Helium Gas Mixtures for Ring Imaging Čerenkov Detectors with CsI Photocathodes, Nucl. Instr. and Meth. A 334, 328 (1993),

http://kirkmcd.princeton.edu/examples/accel/lu_nim_a334_328_93.pdf

11. C. Lu and K.T. McDonald, Properties of Reflective and Semitransparent CsI Photocathodes, Nucl. Instr. and Meth. A 343, 135 (1994),

http://kirkmcd.princeton.edu/examples/accel/lu_nim_a343_135_94.pdf

12. C. Lu, Z. Cheng, K.T. McDonald, D.R. Marlow, E.J. Prebys and R.L. Wixted, Characterization of CsI Photocathodes for Use in a Fast RICH Detector, Nucl. Instr.

and Meth. A 366, 60 (1995), http://kirkmcd.princeton.edu/examples/accel/lu_nim_a366_60_95.pdf

13. C. Lu, Z. Cheng, D.R. Marlow, K.T. McDonald, E.J. Prebys, I.H. Stairs, R.L. Wixted, I. Adachib, R. Itoh, T. Sumiyoshib, N.S. Lockyer and J.E. Millan, Prototype Studies of a Fast RICH Detector with a CsI Photocathode, Nucl. Instr. and Meth. A 371, 155 (1996),

http://kirkmcd.princeton.edu/examples/accel/lu_nim_a371_155_96.pdf

14. (with C. Lu et al.) Aging study for the BESIII-type RPC, Nucl. Instr. and Meth. A

661, S226 (2012), kirkmcd.princeton.edu/examples/detectors/lu_nim_a661_S226_12.pdf

15. R. Acciarri et al., Design and construction of the MicroBooNE detector, Jinst 12,

P02017 (2017), kirkmcd.princeton.edu/examples/detectors/acciarri_jinst_12_p02017_17.pdf

16. M.C. Vignali et al., Deep diffused APDs for charged particle timing applications: Performance after neutron irradiation, Nucl. Instrum. Meth. A 949, 162930 (2020),

http://kirkmcd.princeton.edu/examples/detectors/vignali_nim_a949_162930_20.pdf

17. M.C. Vignali et al., Deep diffused Avalanche photodiodes for charged particles timing, Nucl. Instrum. Meth. A 958, 162405 (2020),

kirkmcd.princeton.edu/examples/detectors/vignali_nim_a958_162405_20.pdf

McDonald Group – Accelerator Physics

1. K.T. McDonald, Fundamental Physics During Violent Accelerations, AIP Conf. Proc.

130, 235 (1985), http://kirkmcd.princeton.edu/examples/accel/mcdonald_aipcp_130_23_85.pdf

2. K.D. Bonin, K.T. McDonald, D.P. Russell and J.8. Flanz, Observation of Interference Between Čerenkov and Synchrotron Radiation, Phys. Rev. Lett. 57, 2264 (1986),

http://kirkmcd.princeton.edu/examples/accel/bonin_prl_57_2264_86.pdf

3. K.T. McDonald, The Hawking-Unruh Temperature and Quantum Fluctuations in Particle Accelerators, Proc. PAC87, p. 1196, kirkmcd.princeton.edu/accel/unruh.pdf

4. K.T. McDonald, Beam Dynamics of the RF Electron Gun of the BNL Accelerator Test Facility, AIP Conf. Proc. 177, 204 (1988),

kirkmcd.princeton.edu/examples/accel/mcdonald_aipcp_177_204_88.pdf

5. K. Batchelor, H. Kirk, J. Sheehan, M. Woodle and K.T. McDonald, Development of a High Brightness Electron Gun for the Accelerator Test Facility at Brookhaven National Laboratory, Proc. EPAC88, p. 954 (1988),

http://kirkmcd.princeton.edu/examples/accel/batchelor_epac88_954.pdf

6. K.T. McDonald, Design of the Laser-Driven RF Electron Gun for the BNL Accelerator Test Facility, IEEE Trans. Electron Devices, 35, 2052 (1988),

http://kirkmcd.princeton.edu/examples/accel/mcdonald_ieeeted_35_2052_88.pdf

7. K.T. McDonald and D.P. Russell, Methods of Emittance Measurement, in Frontiers of Particle Beams: Observation, Diagnosis and Correction, M. Month and S. Turner, eds., (Springer-Verlag, 1989) p. 122,

http://kirkmcd.princeton.edu/examples/accel/mcdonald_emit_meas_89.pdf

8. K.T.McDonald, Beam Dynamics of a Laser-Driven RF Electron Gun, Proc. Linac88 (Williamsburg, Oct. 1988), TH2-04,

http://kirkmcd.princeton.edu/examples/accel/mcdonald_lac88_th2-04.pdf

9. K. Batchelor et al., Operational Status of the Brookhaven National Laboratory Accelerator Test Facility, PAC89, p. 273 (Mar. 20, 1989),

kirkmcd.princeton.edu/examples/accel/batchelor_pac89_273.pdf

10. X.J. Wang, H.G. Kirk, C. Pellegrini, K.T. McDonald amd D.P. Russell, The Brookhaven Accelerator Test Facility Injection System, PAC89, p. 307 (Mar. 20,

1989), kirkmcd.princeton.edu/examples/accel/wang_pac89_307.pdf

11. D.P. Russell and K.T. McDonald, A Beam-Profile Monitor for the BNL Accelerator Test Facility (ATF), PAC89, p. 1510 (Chicago, Mar. 20, 1989),

kirkmcd.princeton.edu/examples/accel/russell_pac89_1510.pdf

12. K. Batchelor et al., Performance of the Brookhaven photocathode rf gun, Nucl. Instr. and Meth. A 318, 92 (1992),

http://kirkmcd.princeton.edu/examples/accel/batchelor_nim_a318_372_92.pdf

13. C. Bula, K.T. McDonald, E. J. Prebys et al., Observation of Nonlinear Effects in Compton Scattering, Phys. Rev. Lett. 76, 3116 (1996),

http://kirkmcd.princeton.edu/examples/QED/bula_prl_76_3116_96.pdf

14. T. Kotseroglou, C. Bula, K.T. McDonald, E. J. Prebys et al., Picosecond Timing of Terawatt Laser Pulses with the SLAC 46-GeV Electron Beam, Nucl. Instr. and Meth.

A 383, 309 (1996), http://kirkmcd.princeton.edu/examples/detectors/kotseroglou_nim_a383_309_96.pdf

15. D.L. Burke, C. Bula, K.T. McDonald, E. J. Prebys et al., Positron Production in Multiphoton Light-by-Light Scattering, Phys. Rev. Lett. 79, 1626 (1997),

http://kirkmcd.princeton.edu/examples/QED/burke_prl_79_1626_97.pdf

16. K.T. McDonald, Comment on “Experimental Observation of Electrons Accelerated in Vacuum to Relativistic Energies by a High-Energy Laser” by Malka et al., Phys. Rev.

Lett. 80, 1350 (1998), http://kirkmcd.princeton.edu/examples/accel/mcdonald_prl_80_1350_98.pdf

17. K.T. McDonald and K. Shmakov, Temporary Acceleration of Electrons While Inside an Intense Electromagnetic Pulse, Phys. Rev. ST Accel. Beams 2, 121301 (1999),

http://kirkmcd.princeton.edu/examples/accel/mcdonald_prstab_2_121301_99.pdf

18. C. Bamber, C. Bula, K.T. McDonald, E. J. Prebys et al., Studies of nonlinear QED in collisions of 46.6 GeV electrons with intense laser pulses, Phys. Rev. D 60, 092004

(1999), http://kirkmcd.princeton.edu/examples/QED/bamber_prd_60_092004_99.pdf

19. K.-J. Kim, K.T. McDonald, G.V. Stupakov and M.S. Zolotorev, Comment on “Coherent Acceleration by Subcycle Laser Pulses”, Phys. Rev. Lett. 84, 3210 (2000),

http://kirkmcd.princeton.edu/examples/accel/kim_prl_84_3210_00.pdf

20. M.V. Berry and K.T. McDonald, Exact and Geometrical-Optics Energy Trajectories in Twisted Beams, J. Opt. A 10, 035005 (2008),

http://kirkmcd.princeton.edu/examples/EM/berry_joa_10_035005_08.pdf

21. G. Alexander et al., Observation of Polarized Positrons from an Undulator-Based Source, Phys. Rev. Lett. 100, 210801 (2008),

kirkmcd.princeton.edu/examples/accel/alexander_prl_100_210801_08.pdf

22. G. Alexander et al., Undulator-based production of polarized positrons, Nucl. Instr. and Meth. A 610, 451 (2009),

http://kirkmcd.princeton.edu/examples/accel/alexander_nim_a610_451_09.pdf

McDonald Group – Muon Collider and Neutrino Factory

An extensive collection of technical notes of this topic, supervised by McDonald is at

http://kirkmcd.princeton.edu/mumu/target/

1. (with C.M. Ankenbrandt et al.) Status of Muon Collider Research and Development and Future Plans, Phys. Rev. ST Accel. Beams 2, 081001 (1999),

http://kirkmcd.princeton.edu/examples/accel/ankenbrandt_prstab_2_081001_99.pdf

2. A. Hassanein et al., An R&D program for targetry and capture at a neutrino factory and muon collider source, Nucl. Instr. and Meth. A503, 70 (2003),

kirkmcd.princeton.edu/examples/accel/hassanenein_nim_a503_70_03.pdf

3. D.B. Cline, J.G. Learned, K. McDonald and F.Sergiampietri, LANNDD – a massive liquid argon detector for proton decay, supernova and solar neutrino studies and a neutrino factory, Nucl. Instr. Meth. A 503, 136 (2003),

http://kirkmcd.princeton.edu/examples/detectors/cline_nim_a503_136_03.pdf

4. M.V. Diwan et al., Very Long Baseline Neutrino Oscillation Experiments for Precise Measurements of Mixing Parameters and CP Violating Effects, Phys. Rev. D 68,

012002 (2003), kirkmcd.princeton.edu/examples/neutrinos/diwan_prd_68_012002_03.pdf

5. M.M. Alsharo’a et al., Status of Neutrino Factory and Muon Collider Research and Development and Future Plans, Phys. Rev. ST Accel. Beams 6, 081001 (2003),

kirkmcd.princeton.edu/examples/accel/alsharo'a_prstab_6_081001_03.pdf

6. K.T. McDonald. A Short Course on Targetry for a Neutrino Factory and Muon Collider (June 4, 2003)

kirkmcd.princeton.edu/mumu/target/targettrans38.pdf

7. N. Simos et al., Target Material Irradiation Studies for High-Intensity Accelerator Beams, Nucl. Phys. B Proc. Suppl. 149, 255 (2005),

kirkmcd.princeton.edu/examples/accel/simos_npb(ps)_149_259_05.pdf

8. V. Graves et al., A free-jet Hg target operating in a high magnetic field intersecting a high-power proton beam, Nucl. Instr. and Meth. A 562, 928 (2006),

kirkmcd.princeton.edu/examples/accel/graves_nim_a562_928_06.pdf

9. N.Simos et al., Solid Target Studies for Muon Colliders and Neutrino Beams, Nucl. Phys. B Proc. Suppl. 155, 288 (2006),

kirkmcd.princeton.edu/examples/accel/simos_npb(ps)_155_288_06.pdf

10. H. Park et al., Optical Diagnostics of Mercury Jet for an Intense Proton Target, Rev. Sci. Instr. 79, 045111 (2008),

http://kirkmcd.princeton.edu/examples/accel/park_rsi_79_045111_08.pdf

11. N. Simos et al., Irradiation damage studies of high power accelerator materials, J.

Nucl. Mat. 377, 41 (2008), http://kirkmcd.princeton.edu/examples/accel/simos_jnm_377_41_08.pdf

12. K.T. McDonald, A Short Course on Targetry for Neutrino Superbeams, Neutrino Factories and Muon Colliders (NuFACT06 Summer Institute, Aug 21, 2006),

kirkmcd.princeton.edu/mumu/target/targettrans53.pdf

13. M. Apollonio et al., Accelerator design concept for future neutrino facilities, J. Inst.

4, p07001 (2009), kirkmcd.princeton.edu/examples/accel/apollonio_jinst_4_p07001_09.pdf

14. R.V. Samulyak, W. Bo, X. Li, H. Kirk and K. McDonald, Computational Algorithms for Multiphase Magnetohydrodynamics and Applications to Accelerator Targets, Cond. Matt. Phys. 13, 43402 (2010),

http://kirkmcd.princeton.edu/examples/accel/samulyak_cmp_13_4_43402_10.pdf

15. D. Adey et al., Light sterile neutrino sensitivity at the nuSTORM facility, Phys. Rev.

D 89, 071301 (2014), kirkmcd.princeton.edu/examples/neutrinos/adey_prd_89_071301_14.pdf

16. Y. Zhan, F. Ladiende, H.G. Kirk and K.T. McDonald, The Effects of Pipe Geometry on Fluid Flow in a Muon Collider Particle Production System, J. Fluid Eng. 136,

101203 (2014), kirkmcd.princeton.edu/examples/fluids/zhan_jfe_136_101203_14.pdf

17. M. Bogomilov et al., Neutrino factory, Phys. Rev. STAB 17, 121002 (2014),

kirkmcd.princeton.edu/examples/accel/bogomilov_prstab_17_121002_14.pdf

18. N. Simos et al., Radiation damage and thermal shock response of carbon-fiber-reinforced materials to intense high-energy proton beams, Phys. Rev. Accel. Beams 19, 111002 (2016),

kirkmcd.princeton.edu/examples/accel/simos_prab_19_111002_16.pdf

19. M. Elbakhshwan et al., X-ray diffraction studies of 145 MeV proton-irradiated AlBeMet 162, Nucl. Mat. Eng. 8, 8 (2016),

http://kirkmcd.princeton.edu/examples/accel/elbakhshwan_nme_8_8_16.pdf

20. N. Simos et al., Multi-MW accelerator target material properties under proton irradiation at Brookhaven National Laboratory linear isotope producer, Phys. Rev. Accel. Beams 21, 053001 (2018),

kirkmcd.princeton.edu/examples/accel/simos_prstab_21_053001_18.pdf

McDonald Group – Daya Bay Reactor Neutrino Experiment

1. L. Ma, C. Lu, K. McDonald et al., The mass production and quality control of RPCs for the Daya Bay experiment, Nucl. Instr. and Meth. A 659, 154 (2011),

kirkmcd.princeton.edu/examples/detectors/ma_nim_a659_154_11.pdf

2. C. Lu, K. McDonald, A.J.S. Smith and J. Zhang, Aging study for the BESIII-type RPC, Nucl. Instr. and Meth. A 661, S226 (2012),

kirkmcd.princeton.edu/examples/detectors/lu_nim_a661_S226_12.pdf

3. J.-L. Xu, C.-G. Lu, K.T. McDonald et al., Design and preliminary test results of Daya Bay RPC modules, Chinese Phys. C 35, 844 (2011),

http://kirkmcd.princeton.edu/examples/detectors/xu_cpc_35_844_11.pdf

4. F.P. An et al., Observation of electron-antineutrino disappearance at Daya Bay, Phys. Rev. Lett. 108, 171803 (2012),

kirkmcd.princeton.edu/examples/neutrinos/an_prl_108_171803_12.pdf

5. F.P. An et al., A side-by-side comparison of Daya Bay antineutrino detectors, Nucl. Instr. and Meth. A 685, 78 (2012),

kirkmcd.princeton.edu/examples/detectors/an_nim_a685_778_12.pdf

6. F.P. An et al., Improved Measurement of Electron Antineutrino Disappearance at Daya Bay, Chinese Phys. C 37, 011001 (2013),

kirkmcd.princeton.edu/examples/neutrinos/an_cp_c37_011001_13.pdf

7. F.P. An et al., Spectral Measurement of Electron Antineutrino Oscillation Amplitude and Frequency at Daya Bay, Phys. Rev. Lett. 112, 061801 (2014),

kirkmcd.princeton.edu/examples/neutrinos/an_prl_112_061801_14.pdf

8. F.P. An et al., Search for a Light Sterile Neutrino at Daya Bay, Phys. Rev. Lett. 113,

141802 (2014), kirkmcd.princeton.edu/examples/neutrinos/an_prl_113_141802_14.pdf

9. F.P. An et al., Independent measurement of the neutrino mixing angle θ_{13} via neutron capture on hydrogen at Daya Bay, Phys. Rev. D 90, 071101 (2014),

kirkmcd.princeton.edu/examples/neutrinos/an_prd_90_071101_14.pdf

10. F.P. An et al., The muon system of the Daya Bay Reactor antineutrino experiment, Nucl. Instr. and Meth. A 773, 8 (2015),

http://kirkmcd.princeton.edu/examples/detectors/an_nim_a773_8_15.pdf

11. F.P. An et al., New Measurement of Antineutrino Oscillation with the Full Detector Configuration at Daya Bay, Phys. Rev. Lett. 115, 111802 (2015),

kirkmcd.princeton.edu/examples/neutrinos/an_prl_115_111802_15.pdf

12. F.P. An et al., The Detector System of the Daya Bay Reactor Neutrino Experiment, Nucl. Instr. and Meth. A 811, 133 (2016),

kirkmcd.princeton.edu/examples/detectors/an_nim_a811_133_16.pdf

13. F.P. An et al., Measurement of the Reactor Antineutrino Flux and Spectrum at Daya Bay, Phys. Rev. Lett. 116, 061801 (2016),

kirkmcd.princeton.edu/examples/neutrinos/an_prl_116_061801_16.pdf

14. F.P. An et al., New measurement of θ_{13} via neutron capture on hydrogen at Daya Bay, Phys. Rev. D 93, 072011 (2016),

kirkmcd.princeton.edu/examples/neutrinos/an_prd_93_072011_16.pdf

15. P. Adamson et al., Limits on Active to Sterile Neutrino Oscillations from Disappearance Searches in the MINOS, Daya Bay, and Bugey-3 Experiments, Phys.

Rev. Lett. 117, 151801 (2016), kirkmcd.princeton.edu/examples/neutrinos/an_prl_117_151801_16.pdf

16. F.P. An et al., Improved Search for a Light Sterile Neutrino with the Full Configuration of the Daya Bay Experiment, Phys. Rev. Lett. 117, 151802 (2016),

kirkmcd.princeton.edu/examples/neutrinos/an_prl_117_151802_16.pdf

17. F.P. An et al., Improved measurement of the reactor antineutrino flux and spectrum at Daya Bay, Chinese Phys. C 41, 013002 (2017),

kirkmcd.princeton.edu/examples/neutrinos/an_cp_c41_013002_17.pdf

18. F.P. An et al., Measurement of electron antineutrino oscillation based on 1230 days of operation of the Daya Bay experiment, Phys. Rev. D 95, 072006 (2017),

kirkmcd.princeton.edu/examples/neutrinos/an_prd_95_072006_17.pdf

19. F.P. An et al., Evolution of the Reactor Antineutrino Flux and Spectrum at Daya Bay, Phys. Rev. Lett. 118, 251801 (2017),

kirkmcd.princeton.edu/examples/neutrinos/an_prl_118_251801_17.pdf

20. F.P. An et al., Seasonal variation of the underground cosmic muon flux observed at Daya Bay, J. Cos. Astro. Phys. 1, 1 (2018),

kirkmcd.princeton.edu/examples/neutrinos/an_jcap01_001_18.pdf

21. F.P. An et al., Cosmogenic neutron production at Daya Bay, Phys. Rev. D 97,

052009 (2018), kirkmcd.princeton.edu/examples/neutrinos/an_prd_97_052009_18.pdf

22. D. Adey et al., Measurement of the Electron Antineutrino Oscillation with 1958 Days of Operation at Daya Bay, Phys. Rev. Lett 121, 241805 (2018),

kirkmcd.princeton.edu/examples/neutrinos/adey_prl_121_241805_18.pdf

23. D. Adey et al., Search for a time-varying electron antineutrino signal at Daya Bay, Phys. Rev. D 98, 0923013 (2018),

kirkmcd.princeton.edu/examples/neutrinos/adey_prd_98_092013_18.pdf

24. D. Adey et al., Improved measurement of the reactor antineutrino flux at Daya Bay, Phys. Rev. D 100, 052004 (2019),

kirkmcd.princeton.edu/examples/neutrinos/adey_prd_100_052004_19.pdf

25. D. Adey et al., A high precision calibration of the nonlinear energy response at Daya Bay, Nucl. Instr. Meth. A 940, 230 (2019),

http://kirkmcd.princeton.edu/examples/detectors/adey_nim_a940_230_19.pdf

26. D. Adey et al., Extraction of the ^{235}U and ^{239}Pu Antineutrino Spectra at Daya Bay, Phys. Rev. Lett. 123, 111801 (2019),

kirkmcd.princeton.edu/examples/neutrinos/adey_prl_121_241805_18.pdf

27. P. Adamson et al., Improved Constraints on Sterile Neutrino Mixing from Disappearance Searches in the MINOS, MINOS+, Daya Bay, and Bugey-3 Experiments, Phys. Rev. Lett. 125, 071801 (2020),

http://kirkmcd.princeton.edu/examples/neutrinos/adamson_prl_125_071801_20.pdf

28. A. Abusleme et al., Optimization of the JUNO liquid scintillator composition using a Daya Bay antineutrino detector, Nucl. Instrum. Meth. A 988, 164823 (2021),

http://kirkmcd.princeton.edu/examples/detectors/abulseme_nim_a988_164823_21.pdf

29. F.P. An et al., Search for electron-antineutrinos associated with gravitational-wave events GW150914, GW151012, GW151226, GW170104, GW170608, GW170814, and GW170817 at Daya Bay, Chinese Phys. C 45, 055001 (2021),

http://kirkmcd.princeton.edu/examples/neutrinos/an_cpc_45_055001_21.pdf

30. F.P. An et al., Antineutrino energy spectrum unfolding based on the Daya Bay measurement and its applications, Chinese Phys. C 45, 073001 (2021),

http://kirkmcd.princeton.edu/examples/neutrinos/an_cpc_45_073001_21.pdf

31. F.P. An et al., Joint Determination of Reactor Antineutrino Spectra from ^{235}U and 239Pu Fission by Daya Bay and PROSPECT, Phys. Rev. Lett. 128, 081801 (2022),

http://kirkmcd.princeton.edu/examples/neutrinos/an_prl_128_081801_22.pdf

32. F.P. An et al., First Measurement of High-Energy Reactor Antineutrinos at Daya Bay, Phys. Rev. Lett. 129, 041801 (2022),

http://kirkmcd.princeton.edu/examples/neutrinos/an_prl_129_041801_22.pdf